P l e a s e h e l p m e s o l v e t h i s
f i n d t h e c a p a c i t y i n l i t r e o f a b u c k e t 2 4 c m i n d i a m e t e r a t t h e t o p , 1 6 c m i n d i a m e t e r a t t h e b o t t o m a n d 1 8 c m d e e p

Answer by rothauserc(4718)

The bucket is a frustum cone and the volume (V) is calculated
V = (pi*h)/3 * (R^2 + r^2 +Rr)
substituting with what we know
V = (3.14*18)/3 * (24^2 +16^2 +24*16)
V = 18.84 * 1216 = 22909.44 cubic centimeters
===============================================================
now we want to find the volume in litres
V = 22909.44 / 1000 = 22.9 litres
================================================================
note that there are a 1000 ml in a litre and a cubic-centimeter = a ml